# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 2.2. Assumptions
The assumptions imposed on the limit random point field are mild. We believe that our mild assumptions allow for the construction of numerous examples of random point fields with Coulomb interactions to which our theorems can be applied.
2.3. Explicit formulae of logarithmic derivatives: Theorem 2.1
Let $\mathbf{I}(\ell+1)$ be as in ((A3)). We define the function $\mathscr{R}_{R,\mathsf{s}}^{\ell}\in C^{1}(\overline{S}_{R})$ by
$\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})=\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}\Big{(}\sum_{s^{i}\not\in\overline{S}_{R}}^{\infty}(-\nabla\Psi)^{(\mathbf{i})}(\xi_{{x}}-s^{i})\Big{)},$ (2.31)
where $\mathsf{s}=\sum_{i}\delta_{s^{i}}$ and $\xi_{{x}}$ is determined via the Taylor expansion of $\nabla\Psi({x}-{y})$ at x=0{x}=0. In (2.31), we take $\xi_{{x}}\in C^{1}(\overline{S}_{R})$ and $\lvert\xi_{{x}}\rvert<\lvert{x}\rvert\leq R$. From (5.19), the infinite sum over ii in (2.31) converges absolutely in $C^{1}(\overline{S}_{R})$ for $\mu$-a.s. $\mathsf{s}$ (see Proposition 5.1).
Let $\eta_{R,\mathsf{s}}^{\mathbf{i}}$ be the constant given by (5.35). Then, $\eta_{R,\mathsf{s}}^{\mathbf{i}}$ and $\mathscr{R}_{R,\mathsf{s}}^{\ell}$ are independent of $\pi_{\overline{S}_{R}}(\mathsf{s})$ such that
$\eta_{R,\mathsf{s}}^{\mathbf{i}}=\eta_{R,\pi_{\overline{S}_{R}^{c}}(\mathsf{s})}^{\mathbf{i}},\quad\mathscr{R}_{R,\mathsf{s}}^{\ell}=\mathscr{R}_{R,\pi_{\overline{S}_{R}^{c}}(\mathsf{s})}^{\ell},$ (2.32)
where the maps $\pi_{\overline{S}_{R}},\pi_{\overline{S}_{R}^{c}}:\mathsf{S}\to\mathsf{S}$ are defined by
$\pi_{\overline{S}_{R}}(\mathsf{s})=\mathsf{s}(\cdot\cap\overline{S}_{R}),\qquad\pi_{\overline{S}_{R}^{c}}(\mathsf{s})=\mathsf{s}(\cdot\cap\overline{S}_{R}^{c}).$
In Theorem 4.1, we establish the existence of the one-point correlation function of $\mu$. Accordingly, the one-reduced Campbell measure is defined by
$\mu^{[1]}(A\times B)=\int_{A}\rho^{1}(x)\,\mu_{x}(B)\,dx,$
where $\mu_{x}=\mu(\cdot+\delta_{x}\mid\mathsf{s}(\{x\})\geq 1)$ denotes the reduced Palm measure of $\mu$ conditioned at $x$.
Theorem 2.1 (Explicit formulae of logarithmic derivatives).
Let $\mathsf{s}=\sum_{i}\delta_{s^{i}}$.
i Assume ((A1))((A3)). Then, $\mu$ possesses a $\bullet$-logarithmic derivative $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$ defined as
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big{(}-\nabla\Phi({x})+\sum_{s^{i}\in\overline{S}_{R}}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}$ $+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big{)}$ (2.33)
for each $R\in\mathbb{N}$ and $\mu^{[1]}$-a.e. $({x},\mathsf{s})\in S_{R}\times\mathsf{S}$, where we set $\mathbb{I}(\ell)=\sqcup_{k=0}^{\ell}\mathbf{I}(k)$. There exists a $\mu^{[1]}$-version of $\mathsf{d}^{\mu}({x},\mathsf{s})$ that is locally Lipschitz continuous in $x$ on
$\mathbb{R}^{d}_{\neq}(\mathsf{s}):=\{x\in\mathbb{R}^{d}:x\neq s^{i}\}.$
ii Assume ((A1))((A3)). Then, for $\mu$-a.s. $\mathsf{s}$ and each $Q\in\mathbb{N}$,
$\lim_{R\to\infty}\eta_{R,\mathsf{s}}^{\mathbf{i}}=0\text{ for all }\mathbf{i}\in\mathbb{I}(\ell),\quad\lim_{R\to\infty}\big{\lVert}\mathscr{R}_{R,\mathsf{s}}^{\ell}\big{\rVert}_{C^{1}(\overline{S}_{Q})}=0,$ (2.34)
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big{(}-\nabla\Phi({x})+\lim_{R\to\infty}\sum_{s^{i}\in\overline{S}_{R}}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}\Big{)}\quad\text{ in $C^{1}(\overline{S}_{Q,\mathsf{s},\epsilon})$}$ (2.35)
for each $\epsilon>0$, where $\overline{S}_{Q,\mathsf{s},\epsilon}=\{{x}\in\overline{S}_{Q};\lvert{x}-s^{i}\rvert\geq\epsilon\text{ for all }i\}$.
iii Assume ((A1))((A4)). Then, the $\bullet$-logarithmic derivative $\mathsf{d}^{\mu}$ of $\mu$ can be expressed as
$\mathsf{d}^{\mu}({x},\mathsf{s})=\lim_{R\to\infty}\beta\sum_{|{x}-s^{i}|\leq R}\frac{{x}-s^{i}}{|{x}-s^{i}\rvert^{d}}.$ (2.36)
The convergence in (2.36) is understood in the sense specified by assumption ((A4)). More precisely, it may hold in one of the following ways: (i) in $L_{\mathrm{loc}}^{2}(\mu^{[1]})$; (ii) $\mu^{[1]}$-almost everywhere; or (iii) for each $\epsilon>0$ and each $Q\in\mathbb{N}$, in $C^{1}(\overline{S}_{Q,\mathsf{s},\epsilon})$ for $\mu$-almost every configuration $\mathsf{s}$, in accordance with (2.30).
Remark 2.6.
Clearly, (2.34) implies that, for $\mu$-a.s. $\mathsf{s}$,
$\lim_{R\to\infty}\beta\Big{(}\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big{)}=0\quad\text{ in $C^{1}(\overline{S}_{Q})$}.$ (2.37)
@@
The existence of the logarithmic derivative $\mathsf{d}^{\mu}$ implies a natural $\mu$-reversible diffusion process [37]. Using $\mathsf{d}^{\mu}$, the ISDE describing the $\mathsf{S}$-labeled dynamics corresponding to the $\mathsf{S}$-valued diffusion process [25] can be written as
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{1}{2}\int_{0}^{t}\mathsf{d}^{\mu}\Big{(}X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}}\Big{)}du,\quad i\in\mathbb{N},$ (2.38)
where $B_{t}^{i}$ are independent standard Brownian motions.
From (2.35), the ISDE (2.38) can be written for each $i\in\mathbb{N}$ as
@@
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}-\frac{\beta}{2}\int_{0}^{t}\nabla\Phi(X_{u}^{i})du+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{|X_{u}^{j}|\leq R,\,j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{d}}du.$
If ((A4)) holds in addition, then, using (2.36), the ISDE (2.38) can be written as
@@
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\lvert X_{u}^{i}-X_{u}^{j}|\leq R,\,j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{d}}du\quad(i\in\mathbb{N}).$
The drift coefficient $(1/2)\mathsf{d}^{\mu}({x},\mathsf{s})$ is not defined on the full configuration space $\mathbb{R}^{d}\times\mathsf{S}$. Consequently, a central difficulty in the analysis of the ISDEs for interacting Brownian motions is to ensure that the particle dynamics $\mathbf{X}$ almost surely avoid the subset where the coefficient is undefined. Moreover, since the gradient of the Coulomb potential $\Psi$, denoted by $\nabla\Psi$, fails to be integrable at infinity, this introduces an additional technical challenge in establishing the existence and uniqueness of solutions. These considerations play a crucial role in rigorously constructing the associated dynamics and in formulating the ISDE in a mathematically precise manner.
@@
Let $\mathfrak{a}\!:\!\mathbb{R}^{d}\!\to\!(\mathbb{R})^{d^{2}}$, where $\mathfrak{a}=(\mathfrak{a}_{kl})_{k,l=1}^{d}$, be a uniformly elliptic, bounded, symmetric matrix-valued function such that $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$. Assume that there exists a constant $c_{2.3}\geq 1$ satisfying
@@
$\nabla\mathfrak{a}=\Big{(}\sum_{k=1}^{d}\frac{\partial\mathfrak{a}_{kl}}{\partial x_{k}}\Big{)}_{l=1}^{d},\quad\mathfrak{b}=\frac{1}{2}\{\nabla\mathfrak{a}+\mathsf{d}^{\mu}\mathfrak{a}\}.$ (2.40)
We generalize ISDE (2.38) to an elliptic-type ISDE such that
$X_{t}^{i}-X_{0}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}(X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}})du\quad(i\in\mathbb{N}).$ (2.41)
Here, $B^{i}=(B^{i,1},\ldots,B^{i,d})$ and $\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}=(\sum_{l=1}^{d}\int_{0}^{t}\sigma_{k,l}(X_{u}^{i})dB_{u}^{i,l})_{k=1}^{d}$.
Loosely speaking, a pair $(\mathbf{X},\mathbf{B})$ of an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued continuous stochastic process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ and an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}=(B^{i})_{i\in\mathbb{N}}$, defined on a filtered space $(\Omega,\mathscr{F},\{\mathscr{F}_{t}\},P)$, is called a weak solution of (2.41) if (2.41) holds for $P$-a.s.
A weak solution $(\mathbf{X},\mathbf{B})$ is called a strong solution if, in addition, $\mathbf{X}$ is measurable with respect to $\mathbf{B}$. See Subsection 13.4 or [34, 16] for the precise definitions.
Furthermore, Subsection 13.4 and [34, 16] also discuss the notions of uniqueness of weak solutions, uniqueness of strong solutions, and pathwise uniqueness.
For $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$, the associated $m$-labeled process $\mathbf{X}^{[m]}$ is given by
@@
Let $\mathfrak{u}\!:\!(\sqcup_{k=0}^{\infty}(\mathbb{R}^{d})^{k})\sqcup(\mathbb{R}^{d})^{\mathbb{N}}\!\to\!\mathscr{M}(\mathbb{R}^{d})$ be the map defined by $\mathfrak{u}((x^{i}))=\sum_{i}\delta_{{x}^{i}}$, where $\mathscr{M}(\mathbb{R}^{d})$ is the set of all measures on $(\mathbb{R}^{d},\mathscr{B}(\mathbb{R}^{d}))$. If $k=0$, we interpret $\mathfrak{u}(x)=\mathsf{0}$ for ${x}\in(\mathbb{R}^{d})^{0}$ by convention, where $\mathsf{0}$ denotes the zero measure.
[…]
$\mathfrak{u}\circ\mathfrak{l}=\mathrm{id.}.$
[…]
From (2.23), $\mathfrak{l}(\mathsf{s})\in(\mathbb{R}^{d})^{\mathbb{N}}$ for $\mu$-a.s. $\mathsf{s}$. For $\mathsf{s}\in\mathsf{S}$, we often write $\mathfrak{l}(\mathsf{s})=(s^{i})_{i}$.
Let $a=\{a(R)\}_{R\in\mathbb{N}}$ be an increasing sequence of natural numbers. Let
$\mathsf{K}[a]=\bigcap_{R=1}^{\infty}\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(S_{R})\leq a(R)\}.$ (2.44)
Then, $\mathsf{K}[a]$ is a compact set in $\mathsf{S}$. Indeed, a subset ${A}\subset\mathsf{S}$ is relatively compact if and only if ${A}\subset\mathsf{K}[a]$ for some $a$ (see [3, 12]). Because $\{\mu^{N_{\mathfrak{n}}}\}_{\mathsf{n}\in\mathbb{N}}$ converges weakly to $\mu$, $\{\mu^{N_{\mathfrak{n}}}\}_{\mathsf{n}\in\mathbb{N}}$ is tight. Hence, for each $\mathsf{r}\in\mathbb{N}$, we can take $a_{\mathsf{r}}=\{a_{\mathsf{r}}(R)\}_{R\in\mathbb{N}}$ such that
$\inf_{\mathfrak{n}\in\mathbb{N}}\mu^{N_{\mathfrak{n}}}(\mathsf{K}[a_{\mathsf{r}}])>1-\frac{1}{\mathsf{r}}.$ (2.45)
From (2.45) and the weak convergence of $\mu^{N_{\mathfrak{n}}}$ to $\mu$, we have
$\mu(\mathsf{K}[a_{\mathsf{r}}])>1-\frac{1}{\mathsf{r}}.$
We note that $\{\mathsf{K}[a_{\mathsf{r}}]\}_{\mathsf{r}\in\mathbb{N}}$ becomes an increasing sequence of compact sets in $\mathsf{S}$.
Let $\mu^{[m]}$ be the $m$-reduced Campbell measure of $\mu$ (cf. [3]). Suppose that the $m$-point correlation function $\rho^{m}$ of $\mu$ exists. Then, we define $\mu^{[m]}$ by
@@
where $\mu_{\mathbf{x}}=\mu(\cdot+\sum_{i=1}^{m}\delta_{{x}^{i}}|\mathsf{s}(\{{x}^{i}\})\geq 1,1\leq i\leq m)$ is the $m$-reduced Palm measure conditioned at $\mathbf{x}=(x^{1},\ldots,x^{m})\in(\mathbb{R}^{d})^{m}$.
Even if $\mu$ does not admit an $m$-point correlation function, we can still define the $m$-reduced Campbell measure by localization. Let $\mu_{\mathsf{r}}=\mu(\cdot\cap K[\mathsf{r}])$. Suppose that $\mu$ has density functions. Then, from (2.44), $\mu_{\mathsf{r}}$ has the $m$-point correlation function $\rho_{\mathsf{r}}^{m}$ and the associated $m$-reduced Campbell measure $\mu^{[m]}_{\mathsf{r}}$ similarly as (2.46). Hence, we define $\mu^{[m]}$ as the increasing limit of $\mu^{[m]}_{\mathsf{r}}$:
$\mu^{[m]}(A)=\lim_{\mathsf{r}\to\infty}\mu^{[m]}_{\mathsf{r}}(A)\quad\text{ for }A\in\mathscr{B}((\mathbb{R}^{d})^{m}\times\mathsf{S}).$ (2.47)
We note that $\mu^{[m]}$ is independent of the particular choice of compact sets $\mathsf{K}[a_{\mathsf{r}}]$. In Theorem 4.1, we will prove that $\mu$ has a $k$-density function on $\overline{S}_{R}^{k}$ for any $R,k\in\mathbb{N}$ and one-point correlation function.
@@
Assume ((A1))((A3)). Then ISDE (2.41) has a weak solution $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$ for $\mu$-a.s. $\mathsf{s}$, such that the associated $m$-labeled process $\mathbf{X}^{[m]}$ is a $\mu^{[m]}$-symmetric and conservative diffusion for each $m\in\{0\}\cup\mathbb{N}$.
We refer to [5] for the concepts of conservative, reversible, and symmetric diffusions. From Theorem 2.1 and Theorem 2.2, we immediately have the following.
Corollary 1.
i Assume ((A1))((A3)). Then, the results in Theorem 2.2 apply to ISDE (2.48):
$X_{t}^{i}-X_{0}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}+\frac{\beta}{2}\int_{0}^{t}\nabla\mathfrak{a}(X_{u}^{i})du$ (2.48)
$+\frac{\beta}{2}\int_{0}^{t}\Big{(}-\nabla\Phi(X_{u}^{i})+\lim_{R\to\infty}\sum_{\lvert X_{u}^{j}\rvert\leq R,\,j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{|X_{u}^{i}-X_{u}^{j}\rvert^{d}}\Big{)}\mathfrak{a}(X_{u}^{i})du\quad(i\in\mathbb{N}).$
ii Assume ((A1))((A4)). Then, the results in Theorem 2.2 apply to ISDE (2.49):
$X_{t}^{i}-X_{0}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})dB_{u}^{i}+\frac{\beta}{2}\int_{0}^{t}\nabla\mathfrak{a}(X_{u}^{i})du$ (2.49)
$+\frac{\beta}{2}\int_{0}^{t}\Big{(}\lim_{R\to\infty}\sum_{\lvert X_{u}^{i}-X_{u}^{j}\rvert\leq R,\,j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{|X_{u}^{i}-X_{u}^{j}\rvert^{d}}\Big{)}\mathfrak{a}(X_{u}^{i})du\quad(i\in\mathbb{N}).$
@@
We next turn to the construction of unique strong solutions to ISDE (2.41). In Theorem 2.3, we establish the existence of such strong solutions and prove their pathwise uniqueness under suitable conditions. To this end, we introduce the notion of the tail decomposition of $\mu$.
Let $\mathscr{T}(\mathsf{S})=\cap_{R=1}^{\infty}\sigma[\pi_{S_{R}^{c}}]$ be the tail $\sigma$-field. Let $\mu_{\mathsf{a}}=\mu(\,\cdot\,|\mathscr{T}(\mathsf{S}))(\mathsf{a})$ be the regular conditional probability of $\mu$ given $\mathscr{T}(\mathsf{S})$. Then,
$\mu=\int\mu_{\mathsf{a}}\mu(d\mathsf{a}).$
Definition 2.3 (Tail decomposition of $\mu$).
A random point field $\mu$ is said to have a tail decomposition if there exists an $\mathsf{S}_{0}\in\mathscr{B}(\mathsf{S})$ such that $\mu(\mathsf{S}_{0})=1$ and, for all $\mathsf{a}\in\mathsf{S}_{0}$,
$\mu_{\mathsf{a}}(\mathsf{A})\in\{0,1\}\text{ for all }\mathsf{A}\in\mathscr{T}(\mathsf{S}),$ (2.50)
@@
This decomposition provides the framework for discussing the construction of strong solutions and the proof of pathwise uniqueness.
@@
i Let $\mathsf{S}_{0}$ be the set in Definition 2.3. Let $\mathsf{a}\in\mathsf{S}_{0}$. Let $\mathsf{A}\in\mathscr{T}(\mathsf{S})$ such that $\mathsf{a}\in\mathsf{A}$. Then, $\mu_{\mathsf{a}}(\mathsf{A})=1$ from (2.51).
ii Similar to [34, Lem. 14.2], we can prove that $\mu$ has a tail decomposition.
Let $\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}$. Let $W(\mathsf{S}_{\mathrm{s,i}})$ be the set of all $\mathsf{S}_{\mathrm{s,i}}$-valued continuous paths defined on $0,\infty)$. We set
$W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})=\{\mathsf{w}\in W(\mathsf{S}_{\mathrm{s,i}});\mathsf{w}_{t}=\sum_{i=1}^{\infty}\delta_{w_{t}^{i}},\mathbf{w}=(w^{i})_{i=1}^{\infty}\in C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}\}.$ (2.53)
For $\mathbf{w}=(w^{i})_{i=1}^{\infty}\in C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$, we set
$I_{R,T}(\mathbf{w})=\sup\{i\in\mathbb{N};\min_{t\in[0,T]}\lvert w^{i}(t)\rvert\leq R\}.$ (2.54)
For a $(\mathbb{R}^{d})^{\mathbb{N}}$-valued process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$, let $\mathsf{X}$ be the associated unlabeled process such that $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$. Note that, if $\mathsf{X}\in W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})$, then $I_{R,T}(\mathbf{X})$ is well defined.
We make the following assumptions. Let $\nu$ be a random point field on $\mathbb{R}^{d}$.

### v4

@@ 2.2. Assumptions
By construction, $\eta_{R,\mathsf{s}}^{\mathbf{i}}$ and $\mathscr{R}_{R,\mathsf{s}}^{\ell}$ are independent of $\pi_{\overline{S}_{R}}(\mathsf{s})$.
@@
The existence of the logarithmic derivative $\mathsf{d}^{\mu}$ yields a natural $\mu$-reversible diffusion [36]. Using $\mathsf{d}^{\mu}$, the ISDE for the labeled dynamics $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ associated with the $\mathsf{S}$-valued diffusion $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$ can be written as in [27]:
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{1}{2}\int_{0}^{t}\mathsf{d}^{\mu}\Big(X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}}\Big)\,du,\quad i\in\mathbb{N}.$ (2.30)
From (2.28), for each $i\in\mathbb{N}$, the ISDE (2.30) can be rewritten as
@@
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}-\frac{\beta}{2}\int_{0}^{t}\nabla\Phi(X_{u}^{i})\,du+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\begin{subarray}{c}\lvert X_{u}^{j}\rvert\leq R\\ j\neq i\end{subarray}}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{d}}\,du,$
and if ((A4)) holds, then (2.29) yields
@@
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\begin{subarray}{c}\lvert X_{u}^{i}-X_{u}^{j}\rvert\leq R\\ j\neq i\end{subarray}}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{d}}\,du,\quad i\in\mathbb{N}.$
Note that the drift coefficient $\frac{1}{2}\mathsf{d}^{\mu}({x},\mathsf{s})$ is not defined on all of $\mathbb{R}^{d}\times\mathsf{S}$. A key point is therefore to ensure that the dynamics $\mathbf{X}$ a.s. avoid the singular set, and to control the long-range divergence coming from the non-integrability of $\nabla\Psi$ at infinity. These issues are central in making the above ISDE formulations rigorous.
@@
In this subsection, we state a generalized ISDE and present the corresponding weak solutions.
Let $\mathfrak{a}\!:\!\mathbb{R}^{d}\!\to\!(\mathbb{R})^{d^{2}}$, where $\mathfrak{a}=(\mathfrak{a}_{kl})_{k,l=1}^{d}$, be a uniformly elliptic, bounded, symmetric matrix-valued function such that $\mathfrak{a}\in C_{b}^{2}(\mathbb{R}^{d})$. Assume that there exists a constant $c_{2.4}\geq 1$ satisfying
@@
$\mathrm{div\,}\mathfrak{a}=\Big(\sum_{k=1}^{d}\frac{\partial\mathfrak{a}_{kl}}{\partial x_{k}}\Big)_{l=1}^{d},\quad\mathfrak{b}=\frac{1}{2}\{\mathrm{div}\mathfrak{a}+\mathsf{d}^{\mu}\mathfrak{a}\}.$ (2.32)
We consider the following elliptic-type ISDE, which generalizes (2.30):
$X_{t}^{i}-X_{0}^{i}=\int_{0}^{t}\sigma(X_{u}^{i})\,dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}\Big(X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}}\Big)\,du,\quad i\in\mathbb{N}.$ (2.33)
Here, $B^{i}=(B_{1}^{i},\ldots,B_{d}^{i})$ is a $d$-dimensional Brownian motion, and
$\int_{0}^{t}\sigma(X_{u}^{i})\,dB_{u}^{i}=\Big(\sum_{l=1}^{d}\int_{0}^{t}\sigma_{k,l}(X_{u}^{i})\,dB_{l}^{i}(u)\Big)_{k=1}^{d}.$
Loosely speaking, a pair $(\mathbf{X},\mathbf{B})$ of an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued continuous process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ and an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued $\{\mathscr{F}_{t}\}$-Brownian motion $\mathbf{B}=(B^{i})_{i\in\mathbb{N}}$, defined on a filtered probability space $(\Omega,\mathscr{F},\{\mathscr{F}_{t}\},P)$, is called a weak solution of (2.33) if (2.33) holds $P$-a.s. A weak solution $(\mathbf{X},\mathbf{B})$ is called a strong solution if, in addition, $\mathbf{X}$ is a measurable function of $\mathbf{B}$. See Subsection 13.5 for precise definitions.
For $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$, the associated $m$-labeled process $\mathbf{X}^{[m]}$ is
@@
Let $\mu^{[m]}$ be the $m$-reduced Campbell measure of $\mu$. If the $m$-point correlation function $\rho^{m}$ of $\mu$ exists, then
@@
where $\mu_{\mathbf{x}}=\mu(\cdot+\sum_{i=1}^{m}\delta_{x^{i}}|\mathsf{s}(\{x^{i}\})\geq 1,\,1\leq i\leq m)$ is the $m$-reduced Palm measure conditioned at $\mathbf{x}=(x^{1},\ldots,x^{m})\in(\mathbb{R}^{d})^{m}$. Even if $\mu$ does not admit an $m$-point correlation function, the $m$-reduced Campbell measure can be defined by localization (see (13.6) and [11]).
@@
Assume ((A1))((A3)). Then ISDE (2.33) has a weak solution $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$ for $\mu$-a.s. $\mathsf{s}$. Moreover, for each $m\in\{0\}\cup\mathbb{N}$, the associated $m$-labeled process $\mathbf{X}^{[m]}$ is a $\mu^{[m]}$-symmetric and conservative diffusion.
@@
We next turn to the construction of unique strong solutions to the ISDE (2.33). In Theorem 2.3, we establish the existence of strong solutions and prove their pathwise uniqueness under suitable conditions. To this end, we introduce the notion of the tail decomposition of the underlying RPF.
Let $\mathscr{T}(\mathsf{S})$ be the tail $\sigma$-field defined by
$\mathscr{T}(\mathsf{S})=\bigcap_{R=1}^{\infty}\sigma[\pi_{S_{R}^{c}}].$
Let $\mu_{\mathsf{a}}=\mu(\,\cdot\mid\mathscr{T}(\mathsf{S}))(\mathsf{a})$ denote the regular conditional probability of $\mu$ given $\mathscr{T}(\mathsf{S})$. Then
$\mu=\int_{\mathsf{S}}\mu_{\mathsf{a}}\,\mu(d\mathsf{a}).$
Definition 2.2 (Tail decomposition).
An RPF $\mu$ is said to admit a tail decomposition if there exists a set $\mathsf{S}_{0}\in\mathscr{B}(\mathsf{S})$ such that $\mu(\mathsf{S}_{0})=1$ and, for all $\mathsf{a}\in\mathsf{S}_{0}$, the following hold:
$\mu_{\mathsf{a}}(\mathsf{A})\in\{0,1\}\quad\text{for all }\mathsf{A}\in\mathscr{T}(\mathsf{S}),$ (2.36)
@@
This decomposition provides a natural framework for the construction of strong solutions and the proof of pathwise uniqueness.
@@
i Let $\mathsf{S}_{0}$ be as in Definition 2.2 and $\mathsf{a}\in\mathsf{S}_{0}$. If $\mathsf{A}\in\mathscr{T}(\mathsf{S})$ satisfies $\mathsf{a}\in\mathsf{A}$, then $\mu_{\mathsf{a}}(\mathsf{A})=1$ by (2.37).
ii As in [34, Lem. 14.2], one can prove that $\mu$ admits a tail decomposition.
Let $W(\mathbb{R}^{d})=C([0,\infty);\mathbb{R}^{d})$. We equip $W(\mathbb{R}^{d})$ with the topology of uniform convergence on compact time intervals and endow $W(\mathbb{R}^{d})^{\mathbb{N}}$ with the corresponding product topology. With this topology, $W(\mathbb{R}^{d})^{\mathbb{N}}$ is a Polish space. Let
$\mathsf{S}_{\mathrm{s,i}}$ $=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\ \mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\},$
$W(\mathsf{S}_{\mathrm{s,i}})$ $=\{\mathsf{w}\,;\,\mathsf{w}\text{ is an }\mathsf{S}_{\mathrm{s,i}}\text{-valued continuous path on }[0,\infty)\}.$
For $\mathbf{w}=(w^{i})_{i=1}^{\infty}\in W(\mathbb{R}^{d})^{\mathbb{N}}$, define the unlabeled path $\mathsf{w}:=\mathfrak{u}_{\mathrm{path}}(\mathbf{w})$ by $\mathsf{w}_{t}=\mathfrak{u}_{\mathrm{path}}(\mathbf{w})_{t}=\sum_{i=1}^{\infty}\delta_{w_{t}^{i}}$. We set
$W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})=W(\mathsf{S}_{\mathrm{s,i}})\cap\mathfrak{u}_{\mathrm{path}}(W(\mathbb{R}^{d})^{\mathbb{N}}),$ (2.39)
$I_{R,T}(\mathbf{w})=\sup\bigl\{i\in\mathbb{N}\,;\,\min_{t\in[0,T]}\lvert w^{i}(t)\rvert\leq R\bigr\}.$ (2.40)
For an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued process $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$, let $\mathsf{X}=\mathfrak{u}_{\mathrm{path}}(\mathbf{X})$ denote the associated unlabeled process. If $\mathsf{X}\in W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})$, then $I_{R,T}(\mathbf{X})$ is well defined.
We impose the following conditions, from basic state-space regularity to absolute continuity with respect to a reference measure $\nu$.